For an ANOVA with three groups that have 10 participants within each, what would be the dfwithin?
Answer #1

Givrn that, for an ANOVA with three groups that have 10 participants within each,

Therefore, number of groups = k = 3

each group has n = 10 participants.

Therefore total sample size = N = nk = 10 * 3 = 30

Degrees of freedom for between groups = k - 1 = 3 -1 = 2

Totak df = N - 1= 30 - 1 = 29

Within df = Total df - between groups df = 29 - 2 = 27

Therefore, dfwithin = 27